A bond is issued with a coupon of 6% paid annually, a maturity of 38 years, and a yield to maturity of 9%. What rate of returURGENT!!
Answer #1

Solution :

It is given that the current maturity = 38 year

YTM = 9%

Face value = 1000

Coupon = 6% = 1000* 6% = 60

f Bond Price = C+1-(1+r) -^ r *(1+r)n

Where r = YTM and n = period

It is given that the investor purchases the bond at 679.28.

Let check the price by using the formula

Price =C+(1-(1+r)-n) + (1 + r)

Price = 60* (1 - (1 + 0.09) -38) 0.09 1000 (1 +0.09)38

0.962174 Price = 60 ** 0.09 1000 26.43668

Price = 641.45 + 26.44 = 679.28

Since investor is selling the bond after one year and YTM is now 12%. The time to maturity is now 37 years.

Let's calculate the price again

Price = 60 * (1-1 +0.12) -37 - 0.12 + 1000 a (1 + 0.12) 37

Price = 60 * 0.984902 1000 0.1266.23184

Price = 492.4508 + 15.09 = 507.55

Price of the bond after one year would be 507.55.

When the investor purchases the bond at 679.28 and holds it for one year then he will receive one coupon payment of $60 as the coupon payment is annually. He will now sell the bond at $507.55

So total loss = selling price - purchase price + coupon received = 507.55 - 679.28 + 60 = -111.73

Rate of return = -111.73 / 679.28 =-16.45%

Answer is -16.45%